# Robust Design and Evaluation of

### v5

@@ Appendix A Proofs of Results in the Main Text
$\mathbb{E}_{n}[(2)]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(3)]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(4)]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(2)]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $00<C_{2}<\infty,C_{3}>$ 0 such that $\mathbb{E}[(3)]<C_{2}$, $\mathbb{E}[(4)]>C_{3}$ under the assumption of bounded nuisance parameter estimators. Putting this together implies that
@@
$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
$\max_{0\leq U\leq 1}(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/3)^{-1}\sum_{i=1}^{n/3}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
@@
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/3}\sum_{i=1}^{n/3}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
Putting together Lemma A.2 and Lemma A.3, the result follows. $\Box$
A.6 Proof of Corollary 4.1
@@
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and $\hat{b}(x)=\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]$.
Under Assumption B.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1. Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
@@
A.7 Proof of Proposition 4.2
In order to prove Proposition 4.2, we first state and prove various auxiliary lemmas about the properties of estimated efficient influence functions. Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
Lemma A.4.
@@
$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\$
Lemma A.5 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
Lemma A.6.
@@
Lemma A.7 (Convergence of plug-in influence function estimator $\phi_{\mu,i}(\hat{\eta})$).
Proof.
This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript $-k$ to ease notation, observe that we can rewrite
Lemma A.8 (Convergence of plug-in influence function estimator $\phi_{\pi\mu,i}(\hat{\eta})$).
Proof.
This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and $-k$. Observe that we can rewrite
$\|\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})\|=$
$+\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
The result then follows. ∎
Lemma A.9.
Proof.
For ease of notation, we omit the dependence on $X_{i}$ and the subscript $-k$. The proof follows an analogous argument to Lemma A.4. Observe that
where
where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma A.6 and Lemma A.7, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma A.6 and Lemma A.8. The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.4 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma A.9. Putting this together, we have shown the first claim
if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
[…]
from which the result follows. $\Box$
Appendix B Oracle inequality for $L_{2}(\mathbb{P})$-error of pseudo-outcome regression
We provide a model-free oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes. This generalizes the pointwise analysis of pseudo-outcome regressions provided in Kennedy(22)-towardsoptimal. In the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
We first state a mild $L_{2}(\mathbb{P})$-stability condition on the researcher’s second-stage nonparametric regression estimator, extending the pointwise stability condition in Kennedy(22)-towardsoptimal.
Assumption B.1.
[…]
We next show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others. This extends Theorem 1 of Kennedy(22)-towardsoptimal, which shows that linear smoothers satisfy an analogous pointwise stability condition.
Proposition B.1.
[…]
Proof.
and this term has mean zero since
by definition of $\widehat{b}$ and iterated expectation. Therefore,
[…]
$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
Further note that $R_{n}^{2}$ equals
[…]
The $L_{2}(\mathbb{P})$-stability condition and the consistency of $\hat{f}(\cdot)$ yields an inequality on the $L_{2}(\mathbb{P})$-convergence of a feasible pseudo-outcome regression relative to an oracle estimator that regresses the true unknown function $f(V_{i})$ on $X_{i}$.
Lemma B.1.
Under the same setup from Assumptions B.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples. Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error. If
  1. i.
  2. ii.
[…]
then
$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
[…]
@@ Appendix B Oracle inequality for $L_{2}(\mathbb{P})$-error of pseudo-outcome regression
$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
Proof.
Note
This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression. In the main text, we apply Lemma B.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
Appendix C Additional Theoretical Results
We collect together additional theoretical results referenced in the main text.
C.1 Linear programming reduction for bounds on positive class performance
As mentioned in Section 4.1 of the main text, the optimization programs defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$, $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$ can be equivalently characterized as linear programs by applying the Charnes-Cooper transformation (CharnesCooper(62)). Lemma C.1 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
Lemma C.1.
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
Proof.
s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$1\displaystyle 0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=$ 1 .
@@ C.2 Variance estimation for overall performance bounds
As mentioned in Section 4.2 of the main text, we now develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance.
Recall from Proposition 4.2, if $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
[…]
@@
[…]
@@
We show that the conditions of Lemma C.2 are satisfied for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$ and $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$. The convergence of probability of the sample estimator $\hat{\Sigma}$ then follows immediately by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.2.
@@
We prove the result for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$ since the analogous argument applies for $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$. Following the proof of Lemma A.7 and Lemma A.8, we observe that
@@
C.3 Estimation under proxy outcome bounds
In this section, we now consider the proxy outcome bounds briefly discussed in Section 2.3.2. Suppose we additionally observe the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to the true outcome $Y_{i}^{*}$ and always observed. As notation, let $\tilde{\mu}_{0}(x)=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ and $\gamma_{1}(x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$. Assume that the proxy outcome satisfies, for all $x\in\mathcal{X}$,

### v6

@@ Appendix B Proofs of Results in the Main Text
$\mathbb{E}_{n}[(\text{\#2})]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(\text{\#3})]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(\text{\#4})]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $0<C_{2}<\infty,C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$, $\mathbb{E}[(\text{\#4})]>C_{3}$ under the assumption of bounded nuisance parameter estimators. Putting this together implies that
@@
$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
@@
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
Putting together Lemma B.8 and Lemma B.9, the result follows. $\Box$
B.7 Proof of Corollary 4.1
@@
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and $\hat{b}(x)=\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]$. Under Assumption A.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1. Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
@@
Appendix C Additional Theoretical Results
C.1 Identification and Estimation of Predictive Disparities
A central question in algorithmic fairness is whether predictive algorithms perform differently across groups defined by sensitive attributes such as race, gender, or income. As noted in Remark 1 of the main text, differences in predictive performance across groups formalize violations of popular fairness criteria. Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$. The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously. In this section, we show that predictive disparities are partially identified under Assumption 2.1. We propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
C.1.1 Overall Predictive Disparities
Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of overall predictive disparities consistent with Assumption 2.1. Further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
Lemma C.1.
where
@@ C.1.1 Overall Predictive Disparities
For $g\in\{0,1\}$ and $\alpha_{g}=P(G_{i}=g)$, observe that
$\mbox{perf}_{g}(s;\beta)=\alpha_{g}^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
We rewrite $\mbox{disp}(s;\beta)$ as
using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it follows that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
The result then follows by noticing that
Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct estimators as in Section 4.1 of the main text. For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for observed outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
Estimation Procedure for Observed Outcome Bounds:
We randomly split the data into $K$ disjoint folds. For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold. For each observation in the $k$-th fold, we construct
for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$. Under regularity conditions and the same arguments as the proof of Proposition 4.1, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
C.1.2 Positive Class Predictive Disparities
We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions. Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Assumption 2.1.
Lemma C.2.
@@ C.1.2 Positive Class Predictive Disparities
Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
@@
We can therefore construct estimators for the non-sharp bounds on positive class disparities by directly applying our estimator in Section 4.2 of the main text. We separately estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$ and then take the appropriate difference. The same arguments as the proof of Proposition 4.2 can be applied to show that this estimator inherits that a partial double robustness property.
@@ C.2 Variance Estimation for Overall Performance Bounds
As mentioned in Section 4.1 of the main text, we develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance under observed outcome bounds. Recall from Proposition 4.1, if $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right),$
@@
To show $\widehat{\Sigma}\xrightarrow{p}\Sigma$, it suffices to show convergence in probability for each entry. We prove this directly by extending Lemma 1 in DornGuoKallus(21).
@@
We show that the conditions of Lemma C.3 are satisfied for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$ and $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$. The convergence of $\hat{\Sigma}$ follows by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.1.
@@
We prove the result for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$ since the analogous argument applies for $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$. Following the proof of Lemma B.4 and Lemma B.5, we observe that
@@
In our Monte Carlo simulations and empirical application, we use this formula for calculating standard errors. In the proof of Proposition 4.1, we show that our estimator is asymptotically equivalent to the centered average of the true bounds, so bootstrap-based inference would also be valid.
C.3 Linear Programming Reduction for Bounds on Positive Class Performance
As mentioned in Section 4.2 of the main text, the optimization program defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$ can be characterized as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)). Lemma C.5 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
Lemma C.5.
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
Proof.
s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
Appendix D Estimation under Alternative Bounding Functions
In the main text, we developed estimation procedures for observed outcome bounds (Section 2.3.1). In this appendix, we extend our estimation framework to two additional classes of bounding functions: proxy outcome bounds (Section 2.3.2) and instrumental variable bounds (Section 2.3.3). Both approaches leverage additional data — proxy outcomes or quasi-experimental variation — to sharpen inference about unobserved confounding. While this complicates notation, we nonetheless can extend our estimators for the bounds on the conditional likelihood and predictive performance measures.
D.1 Estimation under Proxy Outcome Bounds
Suppose the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ is always observed (for both selected and unselected units) and statistically related to the true outcome $Y_{i}^{*}$. We assume that the proxy outcome is equally informative about the true outcome for selected and unselected units: that is, for all for all $x\in\mathcal{X}$,